Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of each bill does Anna have?
Answer by stanbon(75887)

Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of each bill does Anna have?
------------------------
Let # of $5 bills be "x"; value is 5x dollars
Let # of $10 bills be "12-x"; Value is 10(12-x) = 120=10x
------------------
EQUATION:
value + value = 100
5x + 120-10x = 1===============================
Cheers,
Stan H.00
-5x = -20
x = 4 (# of $5 bills)
12-x = 8 (# of $10 bills)
=========================
Cheers,
Stan H.